# Training Neural Networks from Scratch with Parallel Low-Rank Adapters

### v1

@@ 6 Conclusion
In this work, we investigated the feasibility of using low-rank adapters for model pre-training. We introduced LTE, a bi-level optimization method that capitalizes on the memory-efficient properties of LoRA. Although we succeeded in matching performance on moderately sized tasks, several questions remain unresolved. These include: how to accelerate convergence during the final $10\%$ of training; how to dynamically determine the number of ranks or heads required; whether heterogeneous parameterization of LoRA is feasible, where each LoRA head employs a variable rank $r$; and leveraging merging strategies to accompany higher local optimization steps. Our work serves as a proof-of-concept, demonstrating the viability of utilizing low-rank adapters for neural network training from scratch. However, stress tests on larger models are essential for a comprehensive understanding of the method’s scalability. Addressing these open questions will be crucial for understanding the limitations of our approach. We anticipate that our work will pave the way for pre-training models in computationally constrained or low-bandwidth environments, where less capable and low-memory devices can collaboratively train a large model, embodying the concept of the “wisdom of the crowd.
@@ 7 Acknowledgement
JH is supported by the ONR MURI grant N00014-22-1-2740 and the MIT-IBM Watson AI Lab. JB is supported by the MIT-IBM Watson AI Lab and the Packard Fellowship. PI is supported by the Packard Fellowship. BC is supported by the NSF STC award CCF-1231216. We thank Han Guo, Lucy Chai, Wei-Chiu Ma, Eunice Lee, and Yen-Chen Lin for their feedback.
@@ B.1 Is scaling $s$ the same as scaling the learning rate?
$\Delta(\mathbf{B})=-\eta\frac{m_{t}}{\sqrt{v_{t}+\epsilon}}=-\eta\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}s}\hat{m}_{t}}{\sqrt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}s^{2}}\hat{v}_{t}+\epsilon}}=-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}+\epsilon}}$ (7)
However, $\hat{g}_{t}$ is not invariant to $s$. Therefore, while $s$ is not the same as the learning rate, it will impact the downward gradient $\frac{\partial\mathcal{L}(\dots,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}s})}{\partial\mathbf{z}_{out}}$. We will discuss this in the next subsection. It is worth noting that $s$ quadratically impacts the attention map and may require a separate scaling factor as opposed between MLP layers and attention layers. Consider a batched input $\mathbf{X}$ with the output of a linear as $\hat{\mathbf{X}}=\mathbf{W}\mathbf{X}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}s}\mathbf{B}\mathbf{A}\mathbf{X}=\mathbf{W}\mathbf{X}+s\mathbf{D}$. Then, the un-normalized attention map is dominated by the LoRA parameters:
@@
$=\dots+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}s^{2}}\mathbf{W}_{Q}\mathbf{D}\mathbf{D}^{T}\mathbf{W}_{K}^{T}$ (9)

### v2

@@ 6 Conclusion
In this work, we investigated the use of low-rank adapters for model pre-training. We introduced LTE, a bi-level optimization method that capitalizes on the memory-efficient properties of LoRA. Although we succeeded in matching performance on moderately sized tasks, several questions remain unresolved. These include: how to accelerate convergence during the final $10\%$ of training; how to dynamically determine the number of ranks or heads required; whether heterogeneous parameterization of LoRA is feasible, where each LoRA head employs a variable rank $r$; and leveraging merging strategies to accompany higher local optimization steps. Our work serves as a proof-of-concept, demonstrating the viability of utilizing low-rank adapters for neural network training from scratch. However, stress tests on larger models are essential for a comprehensive understanding of the method’s scalability. Addressing these open questions will be crucial for understanding the limitations of our approach. We anticipate that our work will pave the way for pre-training models in computationally constrained or low-bandwidth environments, where less capable and low-memory devices can collaboratively train a large model, embodying the concept of the “wisdom of the crowd.
@@ 7 Acknowledgement
The research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-23-1-0277. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. MH was also supported by the ONR MURI grant N00014-22-1-2740 and the MIT-IBM Watson AI Lab. JB is supported by the MIT-IBM Watson AI Lab and the Packard Fellowship. PI is supported by the Packard Fellowship. BC is supported by the NSF STC award CCF-1231216. We thank Han Guo, Lucy Chai, Wei-Chiu Ma, Eunice Lee, and Yen-Chen Lin for their feedback.
@@ B.1 Is scaling $s$ the same as scaling the learning rate?
$\Delta(\mathbf{B})=-\eta\frac{m_{t}}{\sqrt{v_{t}+\epsilon}}=-\eta\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s}\hat{m}_{t}}{\sqrt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s^{2}}\hat{v}_{t}+\epsilon}}=-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}+\epsilon}}$ (7)
However, $\hat{g}_{t}$ is not invariant to $s$. Therefore, while $s$ is not the same as the learning rate, it will impact the downward gradient $\frac{\partial\mathcal{L}(\dots,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s})}{\partial\mathbf{z}_{out}}$. We will discuss this in the next subsection. It is worth noting that $s$ quadratically impacts the attention map and may require a separate scaling factor as opposed between MLP layers and attention layers. Consider a batched input $\mathbf{X}$ with the output of a linear as $\hat{\mathbf{X}}=\mathbf{W}\mathbf{X}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s}\mathbf{B}\mathbf{A}\mathbf{X}=\mathbf{W}\mathbf{X}+s\mathbf{D}$. Then, the un-normalized attention map is dominated by the LoRA parameters:
@@
$=\dots+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s^{2}}\mathbf{W}_{Q}\mathbf{D}\mathbf{D}^{T}\mathbf{W}_{K}^{T}$ (9)